# Beyond the Silence: How Men Navigate Infertility Through Digital Communities and Data Sharing

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Male infertility accounts for 20–30% of infertility cases (Agarwal et al., 2015), yet remains under-discussed (Patel et al., 2019; Arya and Dibb, 2016), under-supported (Costa Figueiredo et al., 2017; Patel et al., 2019), and heavily stigmatized (Gannon et al., 2004; Dooley et al., 2011). Beyond thwarting reproductive goals, it challenges cultural linkages between masculinity and virility, introducing biological, relational, and psychological uncertainties while offering few social scripts or institutional resources for coping. Unlike many areas of reproductive health that benefit from gendered solidarity and advocacy, men experiencing infertility often face profound isolation. In the absence of formal or culturally accessible support systems, many turn to online spaces for guidance and connection. This study focuses on one such forum, r/maleinfertility described as a space “for men experiencing infertility and male perspectives on infertility. Partners are encouraged to participate, but we ask they post in the daily Partner’s Perspectives thread.”
Research in HCI and CSCW shows how online health and parenting communities foster emotional expression, identity work, and peer support (Ammari et al., 2018; Andalibi and Forte, 2018; Andalibi, 2020). Studies of stay-at-home fathers and non-normative caregivers reveal that men use online forums to reframe vulnerability and caregiving as compatible with strength and responsibility (Ammari et al., 2017, 2019; Lupton et al., 2016), sometimes creating father-only spaces (Ammari and Schoenebeck, 2016). Yet—with the exception of Patel et al. (Patel et al., 2019)—male infertility has received little attention in social computing scholarship, despite its potential to illuminate gendered stigma, identity repair, and support-seeking. Building on this literature, we examine Reddit as a platform where men disclose and process infertility experiences, asking:
@@
While female infertility discourse is well theorized in terms of emotional labor and peer support (Patel et al., 2019), male infertility discourse remains underexplored. Understanding its emotional and linguistic structures can inform platform design and peer support. Prior work suggests that affective and narrative strategies—such as vulnerability or mentoring—boost engagement (Ammari et al., 2025; Andalibi, 2020), but it is unclear which forms of self-expression are most accepted or rewarded in male infertility communities. We therefore also ask:
RQ1b: What types of discourse and user attributes lead to higher engagement in these communities?
Online spaces like Reddit operate across macro (platform-wide), meso (multi-subreddit), and micro (single-subreddit or sub-community) levels of norms (Chandrasekharan et al., 2018; Cai and Wohn, 2021; Ammari et al., 2022; Vachher et al., 2020). These norms shape moderation and acceptable discourse, introducing differences even within one subreddit (Weld et al., 2024). Such dynamics matter in stigmatized contexts, where users may encounter narrow perspectives such as despair narratives or medical fatalism (Randazzo and Ammari, 2025, 2023). To investigate these dynamics, we ask:
RQ2a: Are there distinct sub-communities within male infertility subreddit?
RQ2b: How do different sub-communities frame infertility, masculinity, and coping strategies?
RQ2c: Do these sub-communities function as echo chambers?
Emerging HCI research highlights the role of cross-posters—users who bridge multiple communities—as knowledge brokers and empathy-builders (Ammari et al., 2025). In foster care and trauma recovery spaces, cross-posters act as informal mentors, attract more responses, and introduce diverse viewpoints. In male infertility contexts, cross-posting between male-centric and broader infertility communities (e.g., r/infertility) may broaden resources and normalize emotional expression. Yet there remains more to be known about these dynamics:
@@ 2.1. The Role of Traditional Masculinity Ideologies in Men’s Mental Health
Men’s emotional expression is context-dependent. Limited social settings like funerals or sporting events can permit displays of emotion without threatening masculine norms (Walton et al., 2004), yet pervasive cultural messages such as “boys don’t cry” continue to suppress emotional development (Frosh et al., 2003; Creighton et al., 2013). Dolan et al. (Dolan et al., 2017) found that men facing infertility often avoided sharing their struggles with friends or partners to escape ridicule, adopting a perverse passivity in which infertility was something endured rather than addressed. Our study explores whether digital spaces—offering anonymity and asynchronous interaction—can create alternative contexts that legitimize emotional expression.
@@
Within infertility care, help-seeking is further constrained by the framing of infertility as a female-centered issue, with male discussions focusing narrowly on sperm quality and count (Gannon et al., 2004). Gannon observed that declining sperm counts were depicted as a “war” against men’s bodies, reinforcing narratives of emasculation or quasi-castration. As Patel et al. note (Patel et al., 2019, P.10), such portrayals can deter men from seeking help—whether consulting doctors, confiding in friends, or even talking with partners. If virility and reproduction are central TMI ideals, infertility relegates men, even those otherwise “normal,” to the margins of masculinity.
@@ 2.2. Digital Pathways to Social Connection, Belonging, and Well-being in Men’s Mental Health
Social connectedness is critical to physical and mental health; its absence raises risks of cardiovascular disease and early mortality (Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022). Men may be especially vulnerable because their social networks are often smaller and less emotionally expressive (Steptoe et al., 2013). In this context, digital platforms have emerged as key venues for connection and support in diverse areas including fatherhood (Ammari et al., 2019, 2017), pregnancy loss (Andalibi, 2019), parenting children with special needs (Ammari and Schoenebeck, 2015), and foster care (Lee et al., 2025; Ammari et al., 2025; Lee et al., 2021; Fowler et al., 2022). We focus on two central functions of these communities: (1) building collective knowledge and informal expertise, and (2) (re)building identity and belonging.
@@ 2.2.1. Personal Data Sharing as Collective Knowledge and Informal Expertise
Online support communities allow people to connect with others who share similar experiences (Fox, 2014; Rains and Wright, [n. d.]; Thoits, 2011), offering critical insights and anecdotal evidence to peers and caregivers (Schorch et al., 2016; Soc, 2013). Individuals weigh benefits and drawbacks of data sharing (Wu et al., 2023), but intrinsic rewards such as self-worth and satisfaction motivate ongoing participation (Yan et al., 2016; Soc, 2011).
Self-tracking is common in managing chronic conditions (Li et al., 2010), yet sharing such data can be emotionally fraught (Knittel et al., 2021). Posting personal metrics online provides emotional support, celebrates progress, solicits information, and fosters accountability (Chung et al., 2017; Gui et al., 2017; Lu et al., 2021). These shared narratives help users reinterpret their data, validate experiences, and reinforce coping mechanisms (Knittel et al., 2021). Patel et al. similarly found men sharing detailed fertility regimens and digital narratives of their infertility journeys (Patel et al., 2019, P.9).
For those navigating infertility, chronic illness, or other stigmatized conditions, sharing is more than disclosure: it invites empathy and solidarity (Levonian et al., 2021), normalizes experience, and alleviates shame and uncertainty (Mamykina et al., 2015; Choe et al., 2017). This communal exchange supports emotional healing, motivates continued self-tracking, and co-constructs meaning within online health communities (Jing Wen Lin a, 2023).
@@
A sense of belonging protects against depression and suicidal ideation (McLaren and Challis, 2009), yet strategies for fostering it among socially isolated men are limited. Digital forums connect individuals through shared experience and mutual support (Biagianti et al., 2018). Subjective well-being—spanning emotional stability, social engagement, and personal growth—is similarly vital (Leite et al., 2019), and digital interventions like two-way SMS therapy can rival face-to-face treatment for depression and anxiety (Hull et al., 2020). Effective message design and timing further enhance such interventions (Rahman and Nayan, 2022).
Online health communities extend these benefits by enabling patients and caregivers to provide mutual support (Levonian et al., 2021; Joy et al., 2025), helping men—who often have limited offline networks—combat isolation, find peers, and strengthen relationships (Rains, 2024). For marginalized groups, such communities are essential spaces to organize and receive support, including for stigmatized conditions like obesity (Payne et al., 2025) or mental wellness (Progga et al., 2023). Patel et al. document men explicitly requesting support in these settings (Patel et al., 2019, pp.8-10).
These forums also foster identity reconstruction and psychological resilience, offering safe environments for trauma recovery and emotional validation (Ammari et al., 2025; Randazzo and Ammari, 2025). Designing such spaces requires trauma-informed principles (Randazzo et al., 2023), which emphasize respecting user boundaries and language, avoiding forced disclosure or stigma reproduction, and supporting agency through safe expressive modes like digital storytelling (A. Ankrah et al., 2022). We next discuss key tenets of trauma-informed design.
@@ 2.3. Trauma-informed Design
Trauma-informed design frameworks (Chen et al., 2022; Scott et al., 2023) highlight how social media affordances—such as algorithmic recommendation and comment sorting (Parchoma, 2014)—and governance practices like moderation (Jhaver et al., 2018) can support vulnerable populations. Three core principles guide this work: peer support, user choice, and safety, all underpinned by transparency and trust (Bellini, 2023; Dym and Fiesler, 2020).
@@ 2.3.1. Peer support: Building supportive online spaces
Algorithmic recommendation can enhance transportability, exposing users to narratives that resonate with their experiences (Randazzo and Ammari, 2023; Green, 2021). Yet algorithmic opacity poses risks: emotional content may be amplified to drive engagement, harming well-being and limiting community control (Matias, 2019; Jhaver et al., 2018; Fabbri, 2023).
@@ 2.3.2. Pseudonymity, Algorithmic Curation, and Choice
Reddit’s pseudonymity and limited identity cues facilitate disclosure of stigmatized experiences such as sexual assault, trauma, family struggles, and eating disorders (Andalibi, 2019; Andalibi et al., 2018; Haimson et al., 2015; Randazzo and Ammari, 2023; Chancellor et al., 2018). Users exercise agency by selecting communities (Hwang and Foote, 2021), cross-posting content (Datta et al., 2017), and managing overlapping networks of support. Platform features like badges, awards, and peer recognition further influence participation: awards encourage newcomers, while badges help sustain engagement among committed members (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017).
@@ 2.3.3. Moderation and community governance
Safety depends on effective moderation that defines acceptable discourse, removes harmful content, and bans disruptive users (Jhaver et al., 2018; Mansour and Francke, 2021). Automated tools assist at scale (Gillespie, 2018; Jhaver et al., 2019) but often miss contextual nuance (Ammari et al., 2022). Because definitions of harm are context-specific (Gillespie, 2018, p.98), automated systems struggle with “sticky,” emotionally complex content. Algorithmic opacity can further erode trust when moderation decisions remain unexplained (Kuo et al., 2023).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Male infertility accounts for 20–30% of infertility cases (Agarwal et al., 2015), yet remains under-discussed (Patel et al., 2019; Arya and Dibb, 2016), under-supported (Costa Figueiredo et al., 2017; Patel et al., 2019), and heavily stigmatized (Gannon et al., 2004; Dooley et al., 2011). Beyond thwarting reproductive goals, it challenges cultural linkages between masculinity and virility, introducing biological, relational, and psychological uncertainties while offering few social scripts or institutional resources for coping. Unlike many areas of reproductive health that benefit from gendered solidarity and advocacy, men experiencing infertility often face profound isolation. In the absence of formal or culturally accessible support systems, many turn to online spaces for guidance and connection.
Reddit, a predominantly male platform (Pew Research Center, 2024), has become a critical site for masculine identity (re)construction (LaViolette and Hogan, 2019; Farrell et al., 2019) as socio-economic changes threaten men’s traditionally privileged status (Hanson et al., 2023). This is especially evident in male infertility discussions, where men feel ”a part of their masculinity is being taken away” (Patel et al., 2019), and stigma is so pervasive that men have opted to conduct research interviews in parked cars to consciously ensure their privacy (Dolan et al., 2017). Reddit’s pseudonymity enables engagement with stigmatized topics (Ammari et al., 2019; De Choudhury and De, 2014), fostering solidarity networks (Morini et al., 2025) that protect against depression and suicidal ideation while challenging hegemonic norms of stoicism (Redondo-Sama et al., 2021). The platform’s topic-specific communities (subreddits), moderated by peers facing similar challenges (Gilbert, 2023; Naslund et al., 2016; Proferes et al., 2021) and governed by community-specific rules (Fiesler et al., 2018), enable active support spaces (Morini et al., 2025). This study focuses on r/maleinfertility, described as a space ”for men experiencing infertility and male perspectives on infertility. Partners are encouraged to participate, but we ask they post in the daily Partner’s Perspectives thread.”
HCI and CSCW research shows how online health communities foster emotional expression, identity work, and peer support (Ammari et al., 2018; Andalibi and Forte, 2018; Andalibi, 2020), providing ”social transition machinery” (Haimson, 2018) that helps people understand (Ammari et al., 2025) and heal from trauma (Randazzo and Ammari, 2025). Studies of stay-at-home fathers reveal that men use online forums to reframe vulnerability as compatible with strength (Ammari et al., 2017, 2019; Lupton et al., 2016), sometimes creating father-only spaces (Ammari and Schoenebeck, 2016). Yet—except for Patel et al. (Patel et al., 2019)—male infertility has received little attention in social computing scholarship. We therefore ask:
@@
While female infertility discourse is well theorized (Patel et al., 2019), male infertility remains underexplored. Understanding its emotional and linguistic structures can inform platform design, given that user retention is central to online support community health (Wan et al., 2020; Jo et al., 2023). Prior work suggests affective strategies—sharing vulnerability, mentoring, identity reconstruction—boost engagement (Ammari et al., 2025; Andalibi, 2020; Vermeeren, 2025), but it is unclear which forms of self-expression are most rewarded in male infertility communities. Addressing this lack of clarity regarding male-centric support, we ask:
RQ1b: What types of discourse and user attributes lead to higher engagement in male infertility communities?
Reddit operates across macro, meso, and micro normative levels (Chandrasekharan et al., 2018; Cai and Wohn, 2021; Ammari et al., 2022; Vachher et al., 2020), shaping moderation and acceptable discourse in ways that vary even within subreddits (Weld et al., 2024). Users select communities based on rules (Zou et al., 2024; Proferes et al., 2021), but engage with posts according to personal interest (Yu et al., 2024; Zhang et al., 2017). In stigmatized health contexts, while homophilous groups offer vital support (Randazzo and Ammari, 2023; Randazzo et al., 2023), users may encounter narrow perspectives dominated by despair narratives (Randazzo and Ammari, 2025; Zhong and Gu, 2024; Joy et al., 2025). These dynamics can produce grief bubbles—echo chambers where users risk retraumatization through repeated exposure to similar discussions without encountering alternative perspectives (Randazzo and Ammari, 2025; Goglia and Vega, 2024; Mancini et al., 2022). Drawing on definitions of echo chambers as combining structural closure with topical purity (Morini et al., 2021), we ask:
RQ2a: Are there distinct sub-communities (echo chambers) within the male infertility subreddit?
Online community members adopt distinct functional identities—experts, translators, facilitators (Saxena and Reddy, 2021; Kou et al., 2018). Access to multiple roles proves crucial for identity reconstruction (Laumann, 1973), as individuals navigating life transitions require varied relational support. Ammari et al. (Ammari et al., 2025) demonstrate this multiplicity in foster care communities, where members simultaneously identify as parents, survivors, and professionals—role multiplicity enabling identity reconstruction across temporal stages and relational positions. To examine how this role multiplicity manifests in our context, we ask:
RQ2b: What roles are represented within the different sub-communities in the male infertility subreddit?
The appeal of joining communities stems from discovering one’s ”people” (TeBlunthuis et al., 2022). Homophilous clustering creates protective environments for processing trauma across domains like mental health (Pavalanathan and De Choudhury, 2015), sexual assault (Andalibi et al., 2016), parenting (Ammari et al., 2019, 2017; Ammari and Schoenebeck, 2016), caregiving (Ammari and Schoenebeck, 2015), child loss (Andalibi et al., 2022), and foster care (Ammari et al., 2025). Participation across multiple communities enables tailored self-presentation (Litt and Hargittai, 2016; TeBlunthuis et al., 2022). While cross-posters can be hostile elsewhere (Kumar et al., 2018), in trauma-affected domains they serve as knowledge brokers and empathy-builders (Ammari et al., 2025). In male infertility contexts, cross-posting between male-centric and broader communities may expand resources and normalize emotional expression. To understand the nature of interactions at these community intersections, we ask:
@@ 2.1. The Role of Traditional Masculinity Ideologies in Men’s Mental Health
Men’s emotional expression is context-dependent. Limited social settings like funerals or sporting events can permit displays of emotion without threatening masculine norms (Walton et al., 2004), yet pervasive cultural messages such as “boys don’t cry” continue to suppress emotional development (Frosh et al., 2003; Creighton et al., 2013). Dolan et al. (Dolan et al., 2017) found that men facing infertility often avoided sharing their struggles with friends or partners to escape ridicule, adopting a “perverse passivity” in which infertility was something endured rather than addressed. Our study explores whether digital spaces—offering anonymity and asynchronous interaction—can create alternative contexts that legitimize emotional expression.
@@
Within infertility care, help-seeking is further constrained by the framing of infertility as a female-centered issue, with male-centered discussions focusing narrowly on sperm quality and count (Gannon et al., 2004). Gannon observed that declining sperm counts were depicted as a “war” against men’s bodies, reinforcing narratives of emasculation or quasi-castration. As Patel et al. note (Patel et al., 2019, P.10), such portrayals can deter men from seeking help—whether consulting doctors, confiding in friends, or even talking with partners. If virility and reproduction are central TMI ideals, infertility relegates men, even those otherwise “normal,” to the margins of masculinity.
Traditional masculine ideology—emphasizing self-reliance, emotional restraint, and stoicism—ostracizes men who disclose trauma or support vulnerable community members, creating barriers to help-seeking (Mokhwelepa and Sumbane, 2025). Semaan et al.’s studies of U.S. military veterans demonstrate how ”hyper-masculinity” learned during service impedes post-service disclosure and civilian reintegration (Semaan et al., 2016, 2017), leading veterans to engage in ”hyper-selective disclosure” through information communication technologies (Dosono et al., 2017). These norms shape disclosure even in digital environments, suggesting men with stigmatized health conditions may similarly restrict online help-seeking. Complicating matters further, men perceiving themselves as less masculine report higher harassment endorsement, mediated by masculine norm adherence and toxic disinhibition (Rubin et al., 2020). However, this does not preclude finding support online—communities can support marginalized men, such as transmen navigating identity-related crises (Dym et al., 2019), facilitating identity work difficult in offline contexts constrained by normative expectations.
@@ 2.2. Digital Pathways to Social Connection, Belonging, and Well-being in Men’s Mental Health
Social connectedness is critical to physical and mental health; its absence raises risks of cardiovascular disease and early mortality (Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022). Men may be especially vulnerable because their social networks are often smaller and less emotionally expressive (Steptoe et al., 2013). In this context, digital platforms have emerged as key venues for connection and support in diverse areas including fatherhood (Ammari et al., 2019, 2017; Sepahpour-Fard and Quayle, 2022), pregnancy loss (Andalibi, 2019; Andalibi et al., 2018), parenting children with special needs (Ammari and Schoenebeck, 2015), foster care (Lee et al., 2025; Ammari et al., 2025; Lee et al., 2021; Fowler et al., 2022), and mental health (Morini et al., 2025; De Choudhury et al., 2013; De Choudhury and De, 2014; Saha et al., 2020). We focus on three central functions of these communities: (1) building collective knowledge and informal expertise; (2) care work; and (3) (re)building identity and belonging.
@@ 2.2.1. Personal Data Sharing as Collective Knowledge and Informal Expertise
Online support communities connect people with shared experiences (Fox, 2014; Rains and Wright, [n. d.]; Thoits, 2011), offering critical insights to peers and caregivers (Schorch et al., 2016; Soc, 2013). While individuals weigh benefits and drawbacks of data sharing (Wu et al., 2023), intrinsic rewards like self-worth and satisfaction motivate ongoing participation (Yan et al., 2016; Soc, 2011).
Self-tracking is common in managing chronic conditions (Li et al., 2010), yet sharing such data can be emotionally fraught (Knittel et al., 2021). Through community engagement, posting personal metrics provides emotional support, celebrates progress, solicits information, and fosters accountability (Chung et al., 2017; Gui et al., 2017; Lu et al., 2021), helping users reinterpret data, validate experiences, and reinforce coping mechanisms (Knittel et al., 2021).
Patel et al. similarly found men sharing detailed fertility regimens and digital narratives of their infertility journeys (Patel et al., 2019, P.9). Informational support—knowledge, advice, and guidance—helps people navigate health challenges through peer support offering experiential knowledge (Gage, 2013; MacGeorge et al., 2004) that complements professional care (Levonian et al., 2021). This enables ”collective sensemaking”: collaboratively interpreting health information to co-construct shared understanding (Mamykina et al., 2015; Jing Wen Lin a, 2023) while normalizing experiences, reducing uncertainty, and alleviating shame (Choe et al., 2017).
2.2.2. Care Work in Online Health Communities
MacGeorge et al. (MacGeorge et al., 2004) identified three additional types of support in online health communities. Emotional support offers warmth, empathy, and encouragement through spaces for venting and validation—Progga et al. (Progga et al., 2023) found women with postpartum depression seeking assurance that their emotions were normal, while women experiencing infertility similarly gained support through venting in safe spaces (Zou et al., 2024). Esteem support communicates respect and validation; Payne et al. (Payne et al., 2025) documented fat individuals building solidarity through mutual validation to resist medical stigma, while older members of feminist communities gained standing as veterans with valuable experience (Ammari et al., 2022). Network support cultivates belonging through community ties, reducing isolation and enabling collective action (Ammari and Schoenebeck, 2015)—members support newcomers by inviting direct messages and establishing new social connections (De Choudhury and De, 2014; Ammari et al., 2019, 2025).
@@ 2.2.2. Care Work in Online Health Communities
Network support is vital because belonging protects against depression and suicidal ideation (McLaren and Challis, 2009), yet strategies for cultivating it among isolated men are limited. Digital forums offer connection through shared experience (Biagianti et al., 2018), and well-designed digital interventions can match face-to-face treatment efficacy (Hull et al., 2020; Rahman and Nayan, 2022; Leite et al., 2019).
Online health communities enable patients and caregivers to provide mutual support (Levonian et al., 2021; Joy et al., 2025), helping men—who often lack offline networks—combat isolation (Rains, 2024) while serving as essential organizing spaces for marginalized groups facing stigmatized conditions (Payne et al., 2025; Progga et al., 2023). Men explicitly request support in these settings (Patel et al., 2019, pp.8-10). Yet gender shapes engagement: Zou et al.’s (Zou et al., 2024) analysis of Reddit’s IVF community found women maintaining ”close yet detached” relationships through ”tenuous ties,” with informational support predominating while esteem, network, and tangible support remained rare. Women used the subreddit primarily for venting (emotional support) and information support rather than lasting connections, suggesting online communities may function differently across gendered infertility experiences. Such variation reflects platform affordances and moderation norms. Designing effective spaces requires trauma-informed principles (Randazzo et al., 2023) that respect user boundaries, avoid forced disclosure or stigma reproduction, and support agency through safe expressive modes (A. Ankrah et al., 2022). We discuss these considerations next.
@@ 2.3. Trauma-informed Design
Trauma-informed design frameworks (Chen et al., 2022; Scott et al., 2023) highlight how social media affordances—such as algorithmic recommendation and comment sorting (Parchoma, 2014)—and governance practices like moderation (Jhaver et al., 2018) can support vulnerable populations. Three core principles guide this work: peer support, user choice, and safety, all underpinned by transparency and trust (Bellini, 2023; Dym and Fiesler, 2020). Adhering to these principles fosters identity reconstruction and psychological resilience through safe environments for trauma recovery and emotional validation (Ammari et al., 2025; Randazzo and Ammari, 2025).
@@ 2.3.1. Peer support: Building supportive online spaces
Algorithmic recommendation can enhance transportability, exposing users to narratives that resonate with their experiences which supports peer identification (Randazzo and Ammari, 2023; Green, 2021). Yet algorithmic opacity poses risks: emotional content may be amplified to drive engagement, harming well-being and limiting community control (Matias, 2019; Jhaver et al., 2018; Fabbri, 2023).
@@ 2.3.2. Pseudonymity, Algorithmic Curation, and Choice
Reddit’s pseudonymity and limited identity cues facilitate disclosure of stigmatized experiences such as sexual assault, trauma, family struggles, and eating disorders (Andalibi, 2019; Andalibi et al., 2018; Haimson et al., 2015; Randazzo and Ammari, 2023; Chancellor et al., 2018). Users exercise agency by selecting communities of interest (Hwang and Foote, 2021), cross-posting content (Datta et al., 2017), and managing overlapping networks of support (Ammari et al., 2025). Platform features like badges, awards, and peer recognition further influence participation: awards encourage newcomers, while badges help sustain engagement among committed members (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017) – both of which are important given that retention drives online support community health (Wan et al., 2020; Jo et al., 2023)
@@ 2.3.3. Moderation and community governance
Online community safety depends on moderation that defines acceptable discourse, harmful content, and bans disruptive users (Jhaver et al., 2018; Mansour and Francke, 2021). While automated tools assist at scale (Gillespie, 2018; Jhaver et al., 2019), they miss contextual nuance in emotionally complex content because harm definitions are context-specific (Gillespie, 2018; Ammari et al., 2022, p.98), and algorithmic opacity erodes trust when decisions remain unexplained (Kuo et al., 2023). Human moderators therefore perform essential labor: shaping discussion boundaries, organizing information, applying community norms (Edwards, 2002), and defining membership—as in fathers-only groups excluding mothers (Ammari and Schoenebeck, 2016; Ammari et al., 2019). This constitutes complex care work across multiple power levels, particularly when providing supportive spaces for disclosure while protecting privacy around stigmatizing issues (Saha et al., 2020; Ammari et al., 2022). Drawing on Collins’ matrix of oppression, Gilbert (Gilbert, 2023) shows how r/AskHistorians moderators engage in intensive emotional labor at personal, community, and systemic levels while contending with platform policies and sociotechnical constraints enabling harmful content. This boundary work extends to active advocacy: Wu et al. (Wu et al., 2024) demonstrate how city subreddit moderators establish institutional, cultural, and geographical boundaries to counter racist discourse and brigading. Crucially, moderation functions as both protection and collective empowerment—Ammari et al. (Ammari et al., 2022) show moderators of women-only spaces acting simultaneously as gatekeepers and advocates empowering members to challenge societal norms—revealing moderation as relational and emotional labor that defends boundaries while fostering community agency.
Our study extends this framework to stigmatized populations—men experiencing infertility—examining how moderation tactics protect members from content or actors that could compound medical and emotional distress or violate the trust necessary for vulnerable self-disclosure.